# Privacy-Preserving Collaborative Split Learning Framework for Smart Grid Load Forecasting

### v2

@@ Abstract
Accurate load forecasting is crucial for energy management, infrastructure planning, and demand-supply balancing. The availability of smart meter data has led to the demand for sensor-based load forecasting. Conventional ML allows training a single global model using data from multiple smart meters requiring data transfer to a central server, raising concerns for network requirements, privacy, and security. We propose a split learning-based framework for load forecasting to alleviate this issue. We split a deep neural network model into two parts, one for each Grid Station (GS) responsible for an entire neighbourhood’s smart meters and the other for the Service Provider (SP). Instead of sharing their data, client smart meters use their respective GSs’ model split for forward pass and only share their activations with the GS. Under this framework, each GS is responsible for training a personalized model split for their respective neighbourhoods, whereas the SP can train a single global or personalized model for each GS. Experiments show that the proposed models match or exceed a centrally trained model’s performance and generalize well. Privacy is analyzed by assessing information leakage between data and shared activations of the GS model split.
@@
Split learning, load forecasting, transformers, decentralized learning, privacy preserving, mutual information.
@@ I-A1 Load Forecasting
Figure 1: Recreation of the original Transformer architecture [12].
Following an excellent performance in computer vision (CV) [13] and natural language processing (NLP) community [12], the Transformer architecture [12] (shown in Fig. 1) has recently been employed to capture long-term dependencies in time-series forecasting problems [14, 15, 16]. Instead of working with a single time point at a time (as in RNN), transformer models perform sequence-to-sequence (instead of one sample ahead) forecasting using an encoder-decoder architecture. At the core of transformers, there are self-attention and cross-attention mechanisms which, in vanilla transformer [12], use a point-wise connected matrix leading to a quadratic computational complexity $\mathcal{O}(N^{2})$ w.r.t. the input sequence size.
@@
Method Applied ${}^{\dagger}$ ${}^{\ddagger}$
Tian et al. [9]
SBCTL
S2S RNN
1
Global
$\bigstar$ $\bigstar$ $\bigstar$
Sehova et al. [8]
Central
S2S RNN
1, 24
Global
$\bigstar$ $\bigstar$ $\bigstar$
Fekri et al. [10]
Central
BNLSTM
1 - 200
Global
$\bigstar$ $\bigstar$ $\bigstar$
Yazici et al. [18]
Central
1D-CNN, LSTM, GRU
1, 24
Global
$\bigstar$ $\bigstar$ $\bigstar$
Ryu et al. [11]
Central
MLP-Mixer
24
Global
$\bigstar$ $\bigstar$ $\bigstar$
Taik et al. [19]
FL
LSTM
1
Global
FL
Li et al. [20]
FL
LSTM
1
Global
FL
Liu et al. [21]
FL
LSTM
1, 7
Global
FL + HE - -
Fekri et al. [22]
FL
LSTM
1, 24
Global
FL
Yang et al. [23]
FL
SecureBoost
1
Global
FL
Liu et al. [24]
FL
RNN + LSTM + GRU
0.5 - 4
Global
FL
Husnoo et al. [25]
FL
RNN + LSTM + GRU + CNN
1
Global
FL + DP
Liu et al. [26]
FL
Prob-GBRT
1
Global & Local
FL
Sakuma et al. [27]
SL
S2S LSTM + GRU + 1D-CNN
1, 24
Global & Local
SL
Proposed
SL
S2S FEDformer
96
Neighbourhood
SL + DP
@@ I-A2 Decentralized Learning Methods
Several FL and SL-based methods have been proposed to address these issues. In [19], an FL-based method has been presented for short-term (One-hour) load forecasting for smart meters with similar load profiles. Their approach uses LSTM as the learning model and federated averaging architecture with weighted averaging for model weights aggregation. The method is shown to work well for short-term (one hour ahead) predictions. Similarly, [9] presents a similar approach for short-term forecasting with an emphasis on providing security to the framework via encryption schemes. This, however, leads to increased time complexity of the model. In contrast to [19], authors in [22] compare the performance of two FL techniques, FedSGD (single gradient descent step per client) and FedAVG (multiple gradient updates before merging), and allow their clients to have profiles from different distributions. Similarly, recent studies like [24] and [25] have leveraged RNN, LSTM, and GRU architectures to train global models for short-term forecasting within the FL framework. Furthermore, in the work by [25], differential privacy techniques were applied to obscure signs of shared client gradients, providing an additional layer of protection for client privacy. In a recent work [26], the authors introduce an FL-based boosted multi-task learning framework tailored for inter-district collaborative load forecasting (1 hour ahead). The approach revolves around initially training a central model, which is subsequently employed by individual districts to train personalized models capable of capturing their respective local temporal dynamics. A notable feature of this framework is its use of the probabilistic Gradient-Boosted Regression Tree (GBRT) as the base learner.
In [28], authors use the SL framework to split a 1D CNN network model into two halves and use it to detect heart abnormalities from the medical ECG dataset. Furthermore, they show that in the case of 1D CNN, SL may fail to protect the patients’ private raw data. To mitigate this data leakage, they use differential privacy (DP) [29], where carefully computed noise is added to the patients’ activations as an additional layer of security. Their results show that DP does reduce privacy leakage but at the expense of model performance. Similar to [28], another recent SL-based method [30] splits an LSTM network to train a classifier for time-series data of multiple patients. To reduce privacy leakage, differential privacy has been used to break the 1-1 relationship between input and its split activations.
@@ II-C Differential Privacy
Intuitively, (3) provides an upper bound ($e^{\epsilon}$) on the difference between outputs of the mechanism $\mathcal{M}$ when applied to two neighbouring datasets, where the value of $\epsilon$ controls the overall strength of the privacy mechanism and $\delta$ accounts for the probability that privacy might be violated [32]. Thus, to ensure stronger privacy protection, both $\epsilon$ and $\delta$ should be kept low. With $\delta=0$, the pure $\epsilon-$DP is shown to be much stronger than the $(\epsilon,\delta)-$DP (with $\delta>0$) in terms of mutual information [36]. Nevertheless, the $(\epsilon,\delta)-$DP offers the advantage of advanced composition theorems, enabling a substantially greater number of training iterations compared to pure $\epsilon-$DP with the same $\epsilon$. As a result, most recent works in differentially private machine learning have shifted away from $\epsilon-$DP.
@@ II-D Mutual Information Neural Estimation
The clients’ split layer activations in SL frameworks are shared with the server. However, these activations may carry enough information about the input that an adversary might be able to precisely reconstruct the original data [37]. Various researchers have utilized noise addition mechanisms offered by $\epsilon$-Differential Privacy as a security measure [28] to mitigate this issue. However, it is difficult to quantify the relationship between the added noise level and the information leakage risk. Mutual information (MI) is commonly used in information theory to assess how much information can be inferred from one random variable (RV) about another. Compared to correlations, MI can capture non-linear statistical dependencies between RVs [31]; however, it is difficult to compute, especially for high-dimensional RVs. In [31], the authors have proposed to compute MI using neural networks using the fact that MI between two IID RVs X and Y is equivalent to the Kullback-Leibler Divergence (KLD) between their joint ($\mathbb{P}_{\textbf{XY}}$) and product of their marginal ($\mathbb{P}_{\textbf{X}}\otimes\mathbb{P}_{\textbf{Y}}$) distributions. According to the Donsker-Varadhan representation of KLD [38], the MI between X and Y is lower bound by
@@ III-D Proposed Split Learning Framework
Step 6: Each client computes loss, initiates back-propagation and shares gradients w.r.t. outputs $\textbf{O}_{k,b}^{GS}$ with their GS, that forwards these gradients to SP (lines 20-22 of Algorithm 1).
@@ IV-2 Implementation Detail
TABLE II: Test scores for SplitGlobal (SL-G), SplitPersonal (SL-P), and Central models for different GSs’ neighbourhoods.
@@
SL-G SL-P Central SL-G SL-P Central SL-G SL-P Central
GS 1 0.451 0.453 0.513 0.387 0.383 0.540 0.396 0.405 0.346
GS 2 0.256 0.253 0.246 0.133 0.130 0.133 0.863 0.868 0.855
GS 3 0.372 0.361 0.355 0.256 0.241 0.252 0.663 0.687 0.671
Mean 0.360 0.356 0.371 0.259 0.251 0.308 0.641 0.653 0.624
IV-A Comparison with Centralized Model
In our first experiment, we compare the performance of models trained under the proposed SL framework with a centrally trained FEDformer model. To do so, we train SplitGlobal and SplitPersonal models for 3 GSs and 10 clients per GS. The clients are kept fixed during all training epochs. For the centralized model, we choose the same 10 clients from each GS (30 clients total), as used by our proposed models to train a univariate FEDformer model in a centralized manner. The objective is not only to validate our proposed SL training framework but also to assess the efficacy of training multiple models w.r.t. a single central one. The resulting MAE, MSE, and $R^{2}$ scores on the test sets of clients belonging to different neighbourhoods are summarized in Table II. The MAE, MSE, and $R^{2}$ metric scores for GS 2 and 3 for all 3 models are relatively similar, with the Centrally trained model edging the proposed models slightly in terms of MAE, whereas, in terms of MSE and $R^{2}$, the SplitPersonal model performing better. These results show that all 3 models could generalize the load patterns for GS 2 and 3’s neighbourhood clients. The load profiles of these clients were not too erratic compared to clients belonging to neighbourhood 1, for whom the centrally trained model performed objectively worse.
@@ IV-A Comparison with Centralized Model
When comparing scores of SplitGlobal and SplitPersonal, we see that the latter performs well both in terms of individual GS scores and average ones. This, however, is expected as for SplitPersonal, the SP trains personalized Split-2 networks for each GS. Thus it should be able to generalize well for the respective neighbourhoods. To visualize the model convergence, we present the MSE scores for train, validation, and test sets for the 3 models in Fig. 5. Here, we see that although the central model achieved the lowest training error, its validation and test set errors are the highest, owing to over-fitting to the training data. We can say this because most, if not all, of the difference between central and proposed methods test set scores, is coming from its prediction scores for the GS 1s’ load profiles (see MSE scores given in Table II). As otherwise, the test scores for the central model over GS 1 and 2 were very close to the proposed model’s scores. The run-times of a single epoch for SplitGlobal, SplitPersonal, and Central models are 20, 20, and 40 minutes respectively.
@@ IV-E Privacy preservation using differential privacy
Under the proposed framework, both network splits are trained outside the clients’ premises, and the training is fully controlled by GS and SP entities. Under the honest but curious security assumption, GS and SP shall not deviate from the training process. However, they may try to infer clients’ private data from the received Split-1 activations. In order to secure clients’ private data, we propose to use $\epsilon$-DP [29, 28] to safeguard against such inference attacks. However, before we do this, we first analyze information leakage between the input and output of the proposed Split-1 FEDformer model. To this end, based on the discussion presented in Section II-D, we perform MI estimation using a fully connected neural network (FCNN). A similar strategy has also been used in [42] to infer MI between inputs and gradients of an NN under the FL framework.
@@ IV-E1 Information leakage analysis
Figure 8: The mutual information between input $X_{Inp}$, vs itself, Laplace noise only, clean Split-1 activations $X_{Out}$, and noisy $X_{Out}$ plus Laplace noise $\mathcal{L}(b=2)$. The shown spread signifies one standard deviation.
@@ IV-E2 Analysis under $(\epsilon,\delta=0)-$DP
Figure 9: The mutual information learning trend between input and $\epsilon$-DP protected outputs over a range of $\epsilon$ values.
@@
Figure 10: The prediction test scores and mutual information values for models trained with $\epsilon$-DP over a range of $\epsilon$ values.
@@ IV-E3 Analysis under $(\epsilon,\delta)-$DP
Compared to the stricter pure DP [36], $(\epsilon,\delta)$-DP introduces an extra parameter, $\delta$, to the framework. The rationale behind this enhancement is to provide a level of plausible deniability, allowing for a small probability ($\delta$) that an individual’s data might be exposed or identified. While $\epsilon$ governs the average privacy loss incurred, $\delta$ plays a role in controlling the worst-case privacy loss scenario. Another advantage of $(\epsilon,\delta)-$differential privacy is its advanced composition theorems, which enable significantly more training iterations than pure DP under the same $\epsilon$. This is the reason why most related works in differentially private machine learning have shifted away from pure differential privacy.
@@
To investigate the impact of $(\epsilon,\delta)$-DP on MI leakage and model performance, we repeat the previous experiment but with the $(\epsilon,\delta)$-DP via Gaussian mechanism [40]. In this experiment, we maintained $\epsilon$ at a fixed value of $1.0$ and systematically varied $\delta$ within the range $[0.1,0.05,0.01,0.005,0.001]$, representing a transition from higher worst-case privacy loss to lower levels. The resultant MI leakage, estimated by our estimator (MINE), between clients’ input data and Split-1 layer activations is presented in Figure 11. Interestingly, the variations in $\delta$ had negligible effects on average privacy loss, as $\delta$ primarily governs the worst-case scenario. Furthermore, Figure 12 displays the model’s prediction accuracy over the entire range of $\delta$ values. These metrics indicate that neither MAE nor MSE exhibit substantial variations within the considered range of $\delta$ values. The performance remains consistent and closely mirrors the outcomes observed when $\epsilon=1.0$, as depicted in Figure 10.
@@
IV-F Discussion

### v3

@@ Abstract
Accurate load forecasting is crucial for energy management, infrastructure planning, and demand-supply balancing. The availability of smart meter data has led to the demand for sensor-based load forecasting. Conventional ML allows training a single global model using data from multiple smart meters requiring data transfer to a central server, raising concerns for network requirements, privacy, and security. To alleviate this issue, we propose a split learning-based framework for load forecasting. We split a deep neural network model into two parts, one for each Grid Station (GS) responsible for an entire neighbourhood’s smart meters and the other for the Service Provider (SP). Instead of sharing their data, client smart meters use their respective GSs’ model split for forward passes and only share their activations with the GS. Under this framework, each GS is responsible for training a personalized model split for their respective neighbourhoods, whereas the SP can train a single global or personalized model for each GS. Experiments show that the proposed models match or exceed a centrally trained model’s performance and generalize well. Privacy is analyzed by assessing information leakage between data and shared activations of the GS model split.
@@
Split learning, load forecasting, transformers, decentralized learning, privacy-preserving, mutual information.
@@ I-A1 Load Forecasting
Following an excellent performance in computer vision (CV) [12] and natural language processing (NLP) community [13], the Transformer architecture [13] has recently been employed to capture long-term dependencies in time-series forecasting problems [14, 15, 16]. Instead of working with a single time point at a time (as in RNN), transformer models perform sequence-to-sequence (instead of one sample ahead) forecasting using an encoder-decoder architecture. At the core of transformers, there are self-attention and cross-attention mechanisms which, in vanilla transformer [13], use a point-wise connected matrix leading to a quadratic computational complexity $\mathcal{O}(N^{2})$ w.r.t. the input sequence size.
@@
Method Applied Quantitative Qualitative
Tian et al. [9] SBCTL S2S RNN 1 Global $\bigstar$ $\bigstar$ $\bigstar$
Sehova et al. [8] Central S2S RNN 1, 24 Global $\bigstar$ $\bigstar$ $\bigstar$
Fekri et al. [10] Central BNLSTM 1 - 200 Global $\bigstar$ $\bigstar$ $\bigstar$
Yazici et al. [18] Central 1D-CNN, LSTM, GRU 1, 24 Global $\bigstar$ $\bigstar$ $\bigstar$
Ryu et al. [11] Central MLP-Mixer 24 Global $\bigstar$ $\bigstar$ $\bigstar$
Taik et al. [19] FL LSTM 1 Global FL
Li et al. [20] FL LSTM 1 Global FL
Liu et al. [21] FL LSTM 1, 7 Global FL + HE - -
Fekri et al. [22] FL LSTM 1, 24 Global FL
Yang et al. [23] FL SecureBoost 1 Global FL
Liu et al. [24] FL RNN + LSTM + GRU 0.5 - 4 Global FL
Husnoo et al. [25] FL RNN + LSTM + GRU + CNN 1 Global FL + DP
Liu et al. [26] FL Prob-GBRT 1 Global & Local FL
Sakuma et al. [27] SL S2S LSTM + GRU + 1D-CNN 1, 24 Global & Local SL
Proposed SL S2S FEDformer 96 Neighbourhood SL + DP
@@ I-A2 Decentralized Learning Methods
Several FL and SL-based methods have been proposed to address these issues. In [19], an FL-based method has been presented for short-term (One-hour) load forecasting for smart meters with similar load profiles. Their approach uses LSTM as the learning model and federated averaging architecture with weighted averaging for model weights aggregation. The method is shown to work well for short-term (one hour ahead) predictions. Similarly, [9] presents a similar approach for short-term forecasting with an emphasis on providing security to the framework via encryption schemes. This, however, leads to increased time complexity of the model. In contrast to [19], authors in [22] compare the performance of two FL techniques, FedSGD (single gradient descent step per client) and FedAVG (multiple gradient updates before merging), and allow their clients to have profiles from different distributions. Similarly, recent studies like [24] and [25] have leveraged RNN, LSTM, and GRU architectures to train global models for short-term forecasting within the FL framework. Furthermore, in the work by [25], differential privacy techniques were applied to obscure signs of shared client gradients, providing an additional layer of protection for client privacy. Similarly, in [28], the authors combine FL with the principles of transfer learning to perform demand side forecasting using a Transformer model. In another work [26], the authors introduce an FL-based boosted multi-task learning framework tailored for inter-district collaborative load forecasting (1 hour ahead). The approach revolves around initially training a central model, which is subsequently employed by individual districts to train personalized models capable of capturing their respective local temporal dynamics. A notable feature of this framework is its use of the probabilistic Gradient-Boosted Regression Tree (GBRT) as the base learner.
Moving on to works that leverage for privacy preservation, in [29], authors integrate the FL strategy with local DP to protect user data in recommendation systems. Similarly, in [30], authors enhance user privacy in task-oriented semantic communication within the 6G landscape by incorporating DP and encryption during the training of a deep neural network-based joint source and channel coding (DeepJSCC) model. In [31], authors use the SL framework to split a 1D CNN network model into two halves and use it to detect heart abnormalities from the medical ECG dataset. Furthermore, they show that in the case of 1D CNN, SL may fail to protect the patients’ private raw data. To mitigate this data leakage, they use differential privacy (DP) [32], where carefully computed noise is added to the patients’ activations as an additional layer of security. Their results show that DP does reduce privacy leakage but at the expense of model performance. Similar to [31], another recent SL-based method [33] splits an LSTM network to train a classifier for time-series data of multiple patients. To reduce privacy leakage, differential privacy has been used to break the 1-1 relationship between input and its split activations.
@@ II-C Differential Privacy
Intuitively, (3) provides an upper bound ($e^{\epsilon}$) on the difference between outputs of the mechanism $\mathcal{M}$ when applied to two neighbouring datasets, where the value of $\epsilon$ controls the overall strength of the privacy mechanism and $\delta$ accounts for the probability that privacy might be violated [35]. Thus, to ensure stronger privacy protection, both $\epsilon$ and $\delta$ should be kept low. With $\delta=0$, the pure $\epsilon-$DP is shown to be much stronger than the $(\epsilon,\delta)-$DP (with $\delta>0$) in terms of mutual information [39]. The addition of $\delta$ in the formulation is to provide a level of plausible deniability, allowing for a small probability ($\delta$) that an individual’s data might be exposed or identified by an attacker. While $\epsilon$ governs the average privacy loss incurred, $\delta$ plays a role in controlling the worst-case privacy loss scenario [38]. Additionally, the $(\epsilon,\delta)-$DP offers the advantage of advanced composition theorems, enabling a substantially greater number of training iterations compared to pure $\epsilon-$DP with the same $\epsilon$. As a result, most recent works in differentially private machine learning have shifted away from $\epsilon-$DP.
@@ II-D Mutual Information Neural Estimation
The clients’ split layer activations in SL frameworks are shared with the server. However, these activations may carry enough information about the input that an adversary might be able to precisely reconstruct the original data [40]. Various researchers have utilized noise addition mechanisms offered by $(\epsilon,\delta)$-DP as a security measure to mitigate this issue. However, it is difficult to quantify the relationship between the added noise level and the information leakage risk. Mutual information (MI) is commonly used in information theory to assess how much information can be inferred from one random variable (RV) about another. Compared to correlations, MI can capture non-linear statistical dependencies between RVs [34]; however, it is difficult to compute, especially for high-dimensional RVs. In [34], the authors have proposed to compute MI using neural networks using the fact that MI between two IID RVs X and Y is equivalent to the Kullback-Leibler Divergence (KLD) between their joint ($\mathbb{P}_{\textbf{XY}}$) and product of their marginal ($\mathbb{P}_{\textbf{X}}\otimes\mathbb{P}_{\textbf{Y}}$) distributions. According to the Donsker-Varadhan representation of KLD [41], the MI between X and Y is lower bound by
@@ III-D Proposed Split Learning Framework
Step 6: Each client computes loss, initiates back-propagation and shares gradients w.r.t. outputs $\textbf{O}_{k,b}^{GS}$ with their GS, which forwards these gradients to SP (lines 20-22 of Algorithm 1).
@@ IV-2 Implementation Detail
TABLE II: Test scores for SplitGlobal (SL-G), SplitPersonal (SL-P), Central, and FedAvg models for different GSs’ neighbourhoods.
@@
SL-G SL-P Central FedAvg SL-G SL-P Central FedAvg SL-G SL-P Central FedAvg
GS 1 0.451 0.453 0.513 0.535 0.387 0.383 0.540 0.554 0.396 0.405 0.346 0.342
GS 2 0.256 0.253 0.246 0.268 0.133 0.130 0.133 0.142 0.863 0.868 0.855 0.838
GS 3 0.372 0.361 0.355 0.359 0.256 0.241 0.252 0.273 0.663 0.687 0.671 0.641
Mean 0.360 0.356 0.371 0.387 0.259 0.251 0.308 0.323 0.641 0.653 0.624 0.607
IV-A Comparison with Centralized and FL Models
In our first experiment, we compare the performance of models trained using the proposed SL framework with the FEDformer model trained both centrally and under an FL setting. While FL involves training the entire model on edge clients, in our case, smart meters, this is impractical due to their limited computational capabilities. Nonetheless, we include this comparison for a comprehensive evaluation alongside our approach and the centrally trained model.
To do so, we train SplitGlobal and SplitPersonal models for 3 GSs and 10 clients per GS. The clients are kept fixed during all training epochs. For the centralized model, we choose the same 10 clients from each GS (30 clients total), as used by our proposed models to train a univariate FEDformer model in a centralized manner. The same strategy is used to train a single model in collaboration with 30 clients under FedAvg scheme as well. The objective is not only to validate our proposed SL training framework but also to assess the efficacy of training multiple models w.r.t. the single model under central and FL training strategies. The resulting MAE, MSE, and $R^{2}$ scores on the test sets of clients belonging to different neighbourhoods are summarized in Table II. The MAE, MSE, and $R^{2}$ metric scores for GS 2 and 3 for all 3 models are relatively similar, with the Centrally trained model edging the proposed models slightly in terms of MAE, whereas, in terms of MSE and $R^{2}$, the SplitPersonal model performing better. These results show that all 3 models could generalize the load patterns for GS 2 and 3’s neighbourhood clients. The load profiles of these clients were not too erratic compared to clients belonging to neighbourhood 1, for whom the centrally trained model performed objectively worse.
@@ IV-A Comparison with Centralized and FL Models
When comparing scores of SplitGlobal and SplitPersonal, we see that the latter performs well both in terms of individual GS scores and average ones. This, however, is expected as for SplitPersonal, the SP trains personalized Split-2 networks for each GS. Thus it should be able to generalize well for the respective neighbourhoods. To visualize the model convergence, we present the MSE scores for train, validation, and test sets for the 3 models in Fig. 4. Here, we see that although the central model achieved the lowest training error, its validation and test set errors are the highest, owing to over-fitting to the training data. We can say this because most, if not all, of the difference between central and proposed methods test set scores, is coming from its prediction scores for the GS 1s’ load profiles (see MSE scores given in Table II). As otherwise, the test scores for the central model over GS 1 and 2 were very close to the proposed model’s scores. The run-times of a single epoch for SplitGlobal, SplitPersonal, Central and FL models are 20, 20, 40, and 28 minutes respectively.
@@ IV-E Privacy preservation using differential privacy
Under the proposed framework, both network splits are trained outside the clients’ premises, and the training is fully controlled by GS and SP entities. Under the honest but curious security assumption, GS and SP shall not deviate from the training process. However, they may try to infer clients’ private data from the received Split-1 activations. In order to secure clients’ private data, we propose to use $(\epsilon,\delta)$-DP [32, 38] to safeguard against such inference attacks. However, before we do this, we first analyze information leakage between the input and output of the proposed Split-1 FEDformer model. To this end, based on the discussion presented in Section II-D, we perform MI estimation using a fully connected neural network (FCNN). A similar strategy has also been used in [45] to infer MI between inputs and gradients of an NN under the FL framework.
@@ IV-E1 Information leakage analysis
Figure 7: The mutual information between input $X_{Inp}$, vs itself, Laplace noise only, clean Split-1 activations $X_{Out}$, and $X_{Out}$ plus Laplace noise $\mathcal{L}(b=2)$. The shown spread signifies one standard deviation.
@@ IV-E2 Analysis under $(\epsilon,\delta=0)-$DP
Figure 8: The mutual information learning trend between input and $(\epsilon,\delta=0)$-DP protected outputs over a range of $\epsilon$ values.
@@
Figure 9: The prediction test scores and mutual information values for models trained with $(\epsilon,\delta=0)$-DP over a range of $\epsilon$ values.
@@ IV-E3 Analysis under $(\epsilon,\delta)-$DP
Compared to the stricter pure DP [39], $(\epsilon,\delta)$-DP introduces an extra parameter, $\delta$, into the framework. The reason behind this addition is to provide a level of plausible deniability, allowing for a small probability ($\delta$) that an individual’s data might be exposed or identified by an attacker. While $\epsilon$ governs the average privacy loss incurred, $\delta$ plays a role in controlling the worst-case privacy loss scenario. Another advantage of $(\epsilon,\delta)-$differential privacy is its advanced composition theorems, enabling significantly more training iterations than pure DP under the same $\epsilon$. This is the reason why most related works in differentially private machine learning have shifted away from pure differential privacy.
@@
To investigate the impact of $(\epsilon,\delta)$-DP on MI leakage and model performance, we repeat the previous experiment but with the $(\epsilon,\delta)$-DP via Gaussian mechanism [43]. In this experiment, we maintained $\epsilon$ at a fixed value of $1.0$ and systematically varied $\delta$ within the range $[0.1,0.05,0.01,0.005,0.001]$, representing a transition from higher worst-case privacy loss to lower levels. The resultant MI leakage, estimated by our estimator (MINE), between clients’ input data and Split-1 layer activations is presented in Figure 10. Interestingly, the variations in $\delta$ had negligible effects on average privacy loss, as $\delta$ primarily governs the worst-case scenario. Furthermore, Figure 11 displays the model’s prediction accuracy over the entire range of $\delta$ values. These metrics indicate that neither MAE nor MSE exhibit substantial variations within the considered range of $\delta$ values. The performance closely mirrors the outcomes observed when $\epsilon=1.0$, as depicted in Figure 9.
To illustrate the impact of DP on convergence behavior, Fig. 12 presents the individual loss trends for the training, validation, and test sets over 10 epochs. The plot shows that validation and test losses decrease until about the 5th epoch and then stabilize, while the training loss continues to decline until around the 9th epoch. As expected, the overall loss floor in the DP setting is noticeably higher than in the non-DP case (Fig. 4) due to the added noise. Additionally, training time increased significantly, from roughly $200$ minutes without DP to approximately $454$ minutes with DP, representing a 2.54×2.54$\times$2.54 × increase, primarily due to the computational overhead introduced by noise computations performed through the Diffprivlib library [43].
@@
Figure 12: The convergence graphs for train, validation, and test sets under $(\epsilon=1.0,\delta=0.005)$-DP setting.
IV-F Feature Space Hijacking Attack
Although our framework is built under the ‘honest-but-curious’ assumption, i.e., the utility provider, which is in control of both GS and SP entities, does not intentionally attempt to reconstruct client data or poison the training process, we deliberately relax this assumption to evaluate the worst-case scenario. In this setup, the utility provider (or a compromised entity within it) can deviate from the standard training procedure and attempt to reconstruct private client input data from the shared Split-1 activations using the powerful Feature-Space Hijacking Attack (FSHA) [40].
To compare the effects of DP, we apply the FSHA attack on SplitGlobal model under two scenarios: with and without DP applied to the client side activations. This experiment uses the same settings as used in Section IV-E3. Under this scenario, the smoothed reconstruction MSEs achieved by the FSHA are illustrated in Fig. 13.
  • Without DP, the attack model converges faster and achieves lower reconstruction MSE scores, consistent with observations from previous works [40, 46]. However, even in this case, our proposed framework maintains a reconstruction MSE floor around $0.305$, indicating that the attacker still requires considerable effort to achieve meaningful reconstruction, partly due to the preserved complexity in the input-output relationship at Split-1.
  • With DP, the attack becomes significantly more difficult. The attacker needs a much longer training period, and the resulting reconstruction accuracy is further degraded. This aligns with the intended purpose of DP: to degrade information leakage and protect against reconstruction-based attacks.
The above results are consistent with our analysis under Mutual Information (MI) estimation (section IV-E1) as well. These results demonstrate that our proposed framework, especially when combined with differential privacy, exhibits improved robustness against powerful feature-space attacks.
Figure 13: The reconstruction MSE of the private client data by the FSHA on SplitGlobal model with and without DP.
IV-G Computational and Communication Overhead
In this section, we summarize our findings regarding the various overheads related to our training framework. For details, the reader is referred to the Appendix A.
Communication Overhead
In our framework, the majority of communication overhead is handled between the GS and SP components, with minimal burden placed on the client. For each model update round, the client transmits activations from Split-1, amounting to approximately $6.3$ MB, to the GS and receives a negligible $3$ KB prediction output in return. During backpropagation, the client sends back loss gradients of the same size ($3$ KB), while the heavier $63$ MB gradient tensors are exchanged between the GS and SP. As a result, the client is responsible for only about $4.55$% of the total communication volume, highlighting the communication efficiency of our split learning design. Appendix A-A provides further numerical details.
Computational Overhead
The computational workload in our split learning framework is unevenly distributed between the client (smart meter) and the utility provider (GS-SP). We approximated the number of multiplications for each model component, drawing from the FEDformer architecture. Our analysis shows that the smart meter (Split-1) is responsible for only about $12.66$% of the total forward pass operations, while the remaining 87.34% are executed by the SP (Split-2). This is due to the client’s role being limited to the initial encoding and decoding blocks, while the heavy attention and feed-forward layers are managed by the utility provider. Additionally, the backpropagation and model updates are entirely performed by the GS-SP entities. Since the backpropagation step generally involves around 2.5×2.5$\times$2.5 × (average of 23×2-3$\times$2 - 3 ×) the number of forward-pass operations, the client’s total share of computations across a full model update round amounts to approximately $5$%. This lightweight requirement affirms that smart meters can comfortably participate in training without a significant computational burden. A detailed breakdown is provided in Appendix A-B.
Latency Overhead
Due to the architectural split, most of the computational operations are handled by the GS and SP, which collectively perform over $95$% of the total workload. Consequently, the latency of each training iteration is primarily determined by the utility provider’s infrastructure. Assuming higher-performance hardware at the GS-SP end, the total round-trip latency remains low. While our implementation was done on a single machine, Appendix A-C provides a theoretical latency analysis to offer real-world latency estimates under reasonable hardware assumptions.
Energy Overhead
The client device in our split learning setup is responsible for approximately $5$% of the total computational workload during each model update. This offloading significantly reduces energy consumption at the edge, making the approach attractive for power-constrained devices like smart meters. As shown in Table II, our SplitPersonal model achieved an $18.5$% improvement in prediction performance over the centralized baseline while shifting most of the energy burden to the utility provider’s infrastructure. A deeper analysis of energy distribution and trade-offs is provided in Appendix A-D.
IV-H Discussion